40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 8801     Entry time: Thu Jul 4 01:09:33 2013
Author: Annalisa, Koji 
Type: Update 
Category: LSC 
Subject: PRMI locked with REFL165 

Summary:

- The new REFL165 PD was installed on the AP table
- The REFL165I/Q signals are now showing sensible and robust PRCL/MICH signals
- PRMIsb was locked only with these REFL165 signals

Details:

- Installation of the REFL165 PD

We prepared the REFL165 PD for the 4" optical height. The actual issue was the power supply for the PD.
We soldered wires between the PD and the RF PD interface break-out board. Then the PD interface
cable for the old REFL165 (iLIGO style) was connected.

At the REFL port, most of the light is rejected by the first beam splitter (R=90%?). We attenuated the beam by a factor of 10
using an ND filter. The new PD showed the DC output of ~10V. This corresponds to the photocurrent of 5mA.
(cf. the shot-noise intercept current is ~1mA)

The output of the REFL165 PD was checked with the RF spectrum analyzer. It was a bit surprising but we had a forest of
RF signals betwen 11MHz and 178MHz.
We tried to use a high-pass filter with fc=100MHz (SPH-100) but still the rejection
was not enough. We ended up with using SPH-150 (fc=150MHz).

- Whitening / Demodulation phase

Then we connected the RF output to the SMA cable to the LSC rack. We immediately saw the nice signals from REFL165I/Q
channels, namely sensible structure of pendulum resonances (1/3/16Hz peaks) and floor level.

The whitening level was changed from 21dB to 45dB (max). The DC offsets in the I/Q channels (of the order of 2000~4000)
were removed by the ./LSC/LSCoffset script.

Firstly we locked the PRMI with the usual signals (REFL33I and AS55Q).
The demodulation phase was roughtly tuned (1deg precision) such that the Q phase signal is minimized,
assuming most of the signal is coming from PRCL. Our choise is 74deg.

In this configuration, PRCL shows same quality of signal as our prefered PRCL (i.e. REFL33I) in the amplitude and the sign.

- Locking

We switched to the REFL165 signal by handing off at the input matrix. The input matrix element for REFL165_I was gradually
increasded up to 0.8 while the element for REFL33I was gradually reduced to 0. We did the same for REFL165_Q with the element of 0.2.

Now we tried locking with REFL165I/Q from the beginning. Once the alignment is adjusted, the lock was immediately obtained
only with REFL165I/Q.
Today we did not adjusted the ASC stuff (OPLEVs and PRM ASC) so the lock was not long (<1min). Particularly
ITMX poiting kept drifting and it made the lock difficult. We should check the oplev setup carefully.

- LSC summary

PRCL
Signal source: REFL165I (74deg) / Whitening gain 45dB
Normalization sqrt(POP110I x 0.1) / Trigger POP110I 100up 3down
Servo: input matrix 0.80 -> PRCL Servo FM3/4/5 Always ON G=+2.50
Actuator: output matrix 1.00 -> PRM

MICH
Signal source: REFL165Q (74deg) / Whitening gain 45dB
Normalization sqrt(POP110I x 10.0) / Trigger POP110I 100up 3down
Servo: input matrix 0.20 -> MICH Servo FM4/5 Always On G=-40
Actuator output matrix -1.00 -> ITMX / +1.00 -> ITMY

To Do:

- Refine the PRM asc servo (AC coupled)
- Align oplevs
- ITMX oplev is drifting quickly (~1min time scale)

 

 

 

 

Attachment 1: 130703_PRMI.pdf  17 kB  Uploaded Thu Jul 4 02:49:43 2013  | Hide | Hide all
130703_PRMI.pdf
Attachment 2: Screenshot.png  930 kB  Uploaded Thu Jul 4 02:51:26 2013  | Hide | Hide all
Screenshot.png
ELOG V3.1.3-